OpenStudy (angela210793):

May some1 help me pls?

OpenStudy (anonymous):

yes...

OpenStudy (aroub):

what do u want to know???

OpenStudy (aroub):

mabye i can help

OpenStudy (angela210793):

ME

OpenStudy (angela210793):

Did u see the attachment?

OpenStudy (aroub):

yes

OpenStudy (angela210793):

ME=x=?

OpenStudy (anonymous):

OBF and OAD are similar so OB/Oa=BF/AD=4 ob=4OA OA+OM=AM OB-OM=MB AM=MB OA+OM=OB-OM OA+OM=4OA-OM 2OM=3OA; again ODA &OME are similar ME/AD=OM/OA ME=3/2*AD=3

OpenStudy (anonymous):

angela....is it ok?

OpenStudy (aroub):

i think ur right

OpenStudy (angela210793):

let me see it plsss

OpenStudy (angela210793):

I think ur right 2...Thnx a loooooooooooooot :)

OpenStudy (anonymous):

welcome...